# Evaluating the Performance of ChatGPT for Spam Email Detection

### v1

@@ IV-A Effect of Instance Number in the Prompt
normal-$\bullet$
’Please don’t text me anymore. I have nothing else to say.’ is ”ham”;
$\bullet$ ’The New Jersey Devils and the Detroit Red Wings play Ice Hockey. Correct or Incorrect? End? Reply END SPTV’ is ”spam” .
@@
IV-B Error Analysis
We conduct further research and analysis on the misclassified texts, respectively from the text length and text content comparative study.
TABLE III: Examples of spam texts that incorrectly identified as ham. In the header, “Quant.” stands for “quantitative information”, and “Pub info” for public information.
@@ IV-B Error Analysis
Table III displays the details of eight misclassified texts (they are spam emails but are incorrectly identified as ham). It can be found that the text length of these emails is mostly between 150-160, but the word number distribution has no obvious rules. From the text content analysis, we can find that six of the texts contain quantitative information such as price and phone number. The two pieces of text that did not contain the above information were shorter in length but contained public information such as ”Divorce Barbie” and ”The New Jersey Devils” and were opened as questions, much like small talk between friends, so they were classified as ”ham”.
IV-C Limitations
@@
V Conclusion
This paper explores the capabilities of ChatGPT in spam detection tasks and compares its performance with various baseline models such as SVM, LR, naive Bayes, and BERT. We evaluated the model on two datasets: the Email Spam Detection Dataset and a low-resourced Chinese spam dataset.
From our experimental results demonstrate that ChatGPT performs well in spam classification tasks. However, its overall performance is significantly inferior to supervised models on the English ESD dataset. However, the spam detection ability of ChatGPT is pretty stable across different languages as its performance on the low-resourced Chinese CSD is similchatar to ESD. Another important observation is that ChatGPT has benefits on low-resourced datasets.

### v2

@@ 5.2 Effect of Instance Number in the Prompt
$\bullet$
‘Please don’t text me anymore. I have nothing else to say.’ is “ham”;
$\bullet$ ‘The New Jersey Devils and the Detroit Red Wings play Ice Hockey. Correct or Incorrect? End? Reply END SPTV’ is “spam” .
@@
Table 3: Examples of spam texts that incorrectly identified as ham. In the header, “Quant.” stands for “quantitative information”, and “Pub info” for public information.
@@
5.3 Error Analysis
In order to further elucidate the characteristics of misclassified emails, we conducted an in-depth examination of the textual data. Specifically, we undertook a comparative analysis of the text length and content of the misclassified emails. The results of this analysis are presented in Table 3, which displays the details of eight spam emails that were incorrectly classified as ham. Notably, the text length of these misclassified emails tends to fall within the range of 150-160 words. However, no discernible pattern emerges in terms of word number distribution. A qualitative analysis of the text content reveals that six of the misclassified emails contain quantitative information, including prices and phone numbers. In contrast, the two remaining emails, which were shorter in length, contain references to public information, such as “Divorce Barbie” and “The New Jersey Devils”, and employ a conversational tone reminiscent of small talk between acquaintances. As a result, these emails were mistakenly classified as ham.
5.4 Limitations
@@
6 Conclusion
This paper explores the capabilities of ChatGPT in spam detection tasks and compares its performance with various baseline models such as SVM, LR, naive Bayes, and BERT. We evaluated the model on two datasets: the Email Spam Detection Dataset (ESD) and a low-resourced Chinese spam dataset.
Our experimental results demonstrate that ChatGPT performs well in spam classification tasks, although its overall performance is significantly inferior to supervised models on the English ESD dataset. However, the spam detection ability of ChatGPT is pretty stable across different languages as its performance on the low-resourced Chinese CSD is similar to ESD. Another important observation is that ChatGPT has benefits on low-resourced datasets.
Acknowledgments
The authors thank the reviewers for the valuable comments that helped to improve the paper. This work was supported by “the Fundamental Research Funds for the Central Universities” under Grant 2022114012, and the High-End Foreign Expert Recruitment Program of the Ministry of Science and Technology of the People’s Republic of China, the construction of cutting-edge courses and academic research plans of International Finance and Big Data Management, under Grant G2023138001.